OpenStudy (anonymous):

**Order of Operations** Please help! Will Fan and give Medal!! *Question in Comments*

OpenStudy (anonymous):

Here it is

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

@Candy13106 @alondrachaire @Objection @purpleluffa @TheWaffleMan149 Please help me!

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

All of them

OpenStudy (anonymous):

Well the first one is 49

OpenStudy (anonymous):

Because -7 x -7 is 49

OpenStudy (anonymous):

Ohh. Whats the second?

OpenStudy (anonymous):

4^3 is the same thing as 4 x 4 x 4. 4x4=16. 16x4 is 64

OpenStudy (anonymous):

so its 64

OpenStudy (anonymous):

Okay! What about the rest? thanks alot to! :)

OpenStudy (anonymous):

No problem! 2^4 is the same thing as 2x2x2x2. 2x2 is 4. 4x2=8. 8x2=16, so 3 is 16

OpenStudy (anonymous):

Now for number 4, its a whole equation.

OpenStudy (anonymous):

And the order of operations goes PEMDAS (parentheses, then exponents, then multiplication/division, then addition/subtraction)

OpenStudy (anonymous):

So exponents first

OpenStudy (anonymous):

So 4^2=16

OpenStudy (anonymous):

I'm sorry, I'd help you but I'm literally so terrible at math, that's why I'm here, haha. But I can totally help you with literature and other things :)

OpenStudy (anonymous):

16+7=23. 23-2 =21, so its 21

OpenStudy (anonymous):

@purpleluffa Lol I gotcha! Thanks a lot though! :)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

For the last one, the parentheses are first, so you do 6x0.5 first, which is 3

OpenStudy (anonymous):

Then you do the exponents, which is 3^2, which is 9

OpenStudy (anonymous):

Then the multiplication, which is now 6x9, which is 54. Lastly, you do 54+3, which is 57. Thats the answer!

OpenStudy (anonymous):

Sweet! Thanks man!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

@TheWaffleMan149 Are you good with health education? If so i have some questions with this one

OpenStudy (anonymous):

Well I don't know a thing about that kind of stuff, but common sense tells me that eating junk food and playing computer games will most likely lead to illness in that situation

OpenStudy (anonymous):

4 is setting goals

OpenStudy (anonymous):

5 is advocacy

OpenStudy (anonymous):

I am FAIRLY SURE that 3 is gaining awareness

OpenStudy (anonymous):

And thats all of them :)

OpenStudy (anonymous):

@TheWaffleMan149 Thanks man!

OpenStudy (anonymous):

Any time!

OpenStudy (anonymous):

@TheWaffleMan149 you know dis?

OpenStudy (anonymous):

The one about the mayans

OpenStudy (anonymous):

Yooo!! @TheWaffleMan149 I got 100% on that one!

OpenStudy (anonymous):

Glad I could help!

OpenStudy (anonymous):

:D